OpenStudy (anonymous):

y = -x + 3 y = x - 3 what is the solution to the system of equations?

OpenStudy (tommo_lcfc):

Equate the terms so -x + 3 = x - 3 -2x = -6 so x = 3.

sam (.sam.):

y=0

Parth (parthkohli):

Adding both equations- \[2y = 0\] x = y + 3 x = 3. So, \[y = 0\] \[x = 3\]

OpenStudy (tommo_lcfc):

Yes and that! Thanks @.Sam and @ParthKohli.

OpenStudy (anonymous):

is it (-3,0)